OpenStudy (anonymous):

The area of a rectangle is 42 yd ^2 , and the length of the rectangle is 5 yd more than twice the width. Find the dimensions of the rectangle.

OpenStudy (anonymous):

Hello Paullee I can help you with this. This is one of my favorite types of math problems

OpenStudy (anonymous):

k thx appreciate it because I'm stuck

OpenStudy (anonymous):

Just a sec, I am having problems with this one too lol, I know the answer is 12 and 3.5 but trying to get the work to be typed correctly

OpenStudy (anonymous):

So, you have area = 42 and your length is 5 longer than 2widths L*w=A L+5=2w (2w-5)w=42 2w^2-5w-42=0 (w-12)(w+7)=0, w has to be greater than 0 w=12 L*w=42 L*12=42 L=3.5 I think that is right :P

OpenStudy (anonymous):

so: L=3.5 & W=12?

OpenStudy (anonymous):

I think it should be the other way around actually O.O oh duh, the original second equation should be w+5=2L, sorry about that :P so just switch the two terms in everything

OpenStudy (anonymous):

So L=12 & W=3.5

OpenStudy (anonymous):

because this the example they gave me looking like this to fill in Length: Width:

OpenStudy (anonymous):

Well 12 times 3.5 is 42, and and 12 is 5 more than double 3.5

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Np bro

OpenStudy (anonymous):

it was right

OpenStudy (anonymous):

your the best